OpenStudy (anonymous):

Solve the given system, using the substitution method. y = 3x – 4 9x – 3y = 14

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

A. There is no solution. B. (12, 14) C. (14, 12) D. There are an infinite number of solutions.

hartnn (hartnn):

plug in y= 3x -4 in the 2nd equation 9x - 3(3x-4) = 14 can you find 'x' from here ? try

OpenStudy (anonymous):

I tried that and got \[9x-9x-12=14\]

hartnn (hartnn):

-3*-4 is actually +12

OpenStudy (anonymous):

-.- I feel stupid :P

hartnn (hartnn):

and note that 9x-9x =0

hartnn (hartnn):

so, 12 = 14 is that even possible ?

OpenStudy (anonymous):

So, its A no solution

hartnn (hartnn):

absolutely correct :)

OpenStudy (anonymous):

Thanks for helping me do it myself :)

hartnn (hartnn):

most welcome ^_^

OpenStudy (anonymous):

Instead of giving the answer you helped me understand